OpenStudy (anonymous):

Derive f ' (x) for f(x) = 2x^2+2x/(1+2x)^2

OpenStudy (anonymous):

\[2x ^{2}+2x \div(1+2x)^{2}\]

OpenStudy (anonymous):

use the difference rule. \[(\frac{ u }{v })' = \frac{ \left( (u' * v) + v' * u\right) }{ v ^{2} }\]

OpenStudy (anonymous):

In other words, let the numerator equal to u and the denominator equal to v and then solve.

OpenStudy (anonymous):

I did. I'm having trouble simplifying it. Here's where I am stuck \[(1+2x)^{2}(2)(2x+1)-2x(x+1)2(1+2x)(2)\div(1+2x)^{4}\]

OpenStudy (anonymous):

do you know chain rule?

OpenStudy (anonymous):

you can avoid using quotiont rule and turn it into a product rule problem as: \[f(x)=2x^2+2x(1+2x)^{-2}\]

OpenStudy (anonymous):

interesting... I do know the chain rule. I'm having trouble simplifying it though.

OpenStudy (anonymous):

(1+2x)2(2)(2x+1)−2x(x+1)2(1+2x)(2)÷(1+2x)4 Once I'm here I don't know where to go next.

OpenStudy (anonymous):

there is always more than 1 way to eat a pie

OpenStudy (anonymous):

especially in math

OpenStudy (anonymous):

\[ f'(x)=4x+\left[(2)(1+2x)^{-2}+(2x)(-2)(1+2x)^{-3}(2)\right] \] simplify and thats the answer using my method!

OpenStudy (anonymous):

can you re-write your step legibly?

OpenStudy (anonymous):

I did but I think its wrong... Let me show you...

OpenStudy (anonymous):

\[4\div(1+2x)^{5}\]

OpenStudy (anonymous):

that small? here using quotient rule: \[ f'(x)=4x+\frac{(2)(1+2x)^2-(2x)(2)(1+2x)^1(2)}{(1+2x)^4} \]

OpenStudy (anonymous):

is it 4/(1+2x)^3?

OpenStudy (anonymous):

how? \[ f'(x)=4x+\frac{2(1+2x)-(2x)(4)}{(1+2x)^3}\\ f'(x)=4x+\frac{2(1-2x)}{(1+2x)^3} \] i do not think this can be simplified further

OpenStudy (anonymous):

what happens to the (2x)(4)?

OpenStudy (anonymous):

the numerator of the second term got simplified